OpenStudy (anonymous):

A manufacturer produces soda cans and a quality control worker randomly selects two cans from the assembly line for testing. Past statistics show that 14% of the cans are defective. What is the probability that the two selected cans are defective if the quality control worker selects the two cans from a batch of 50 cans? P(Both defective) = 7/50 P(Both defective) = 49/2500 P(Both defective) = 3/175 P(Both defective) = 7/25

OpenStudy (aveline):

Past statistics show that 14% of the cans are defective. Two cans are selected from a batch of 50 cans. What is 14% of 50?

OpenStudy (aveline):

@lydiasantos

OpenStudy (anonymous):

THanks Aveline, in that case its 7/50 ??

OpenStudy (anonymous):

because 14% of 50 is 7

OpenStudy (anonymous):

right?

OpenStudy (aveline):

Actually 7/50 is the chance of selecting one defective can. If we want to find the chance of selecting two defective cans, we need to multiply. \[\frac{ 7 }{ 50 }\times \frac{ 7 }{ 50 }\]

OpenStudy (anonymous):

oh okay, so then its 49/2500 ??

OpenStudy (aveline):

Yes ^^

OpenStudy (anonymous):

cool thanks